# Non-Stationary Lipschitz Bandits

### v1

@@ Appendix D Proof of the upper bound of MDBE (Theorem 2)
$\textsc{Replay}(s,d)$ denotes a replay at depth $d$ starting at round $s$.
@@ D.1 Bias of the mean estimates
$\forall[s_{1},s_{2}],\,\forall d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t},\,\forall B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]},\quad\left|\sum_{t=s_{1}}^{s_{2}}\mathbb{E}_{t-1}\left[\hat{\delta}_{t}(B^{\prime},B)\right]-\sum_{t=s_{1}}^{s_{2}}\delta_{t}(B^{\prime},B)\right|\leq\frac{4(s_{2}-s_{1})}{2^{d}}\,.$
Proof of Proposition 3.
@@
From Assumption 1, for all bin $B\in\mathcal{T}_{d}$ and any $x\in B$, we have:
@@
Summing this inequality over the interval $[s_{1},s_{2}]$ establishes the second claim. ∎
@@ D.2 Concentration of the mean estimates
Our sampling scheme (Algorithm 2) is carefully constructed to ensure that the estimated gap between any two active bins at any active depth concentrates around its conditional expectation, especially by controlling its cumulative variance. Our result relies on this following inequality.
Lemma 1 (Beygelzimer et al. (2011)).
@@
Let $\mathcal{E}_{1}$ be the following event: for all intervals $[s_{1},s_{2}]$, depths $d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t}$, and bins $B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$, we have
@@
$\geq\frac{1}{|\mathcal{B}_{t}(d^{\prime})\cap\mathrm{Children}(B_{d},d^{\prime})|}\frac{1}{2^{d}}\geq\frac{2^{d}}{2^{d^{\prime}}}\frac{1}{2^{d}}=\frac{1}{2^{d^{\prime}}}.$
Bound on each $M_{t}$. Using the definition of $\hat{\delta}_{t}(B^{\prime},B)$ from (2),
@@
$\mathbb{E}_{t-1}\left[\hat{\delta}_{t}^{2}(B^{\prime},B)\right]\leq\frac{2}{(1/2^{d})^{2}}=2^{d+1}.$
Final concentration bound. Applying Lemma 1 over any interval $[s_{1},s_{2}]$ for fixed $d$ and bins $B,B^{\prime}$, we get that with probability at least $1-\delta$,
@@
Choosing $\delta=1/T^{7}$ and ${\color[rgb]{0.68,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0,0}\pgfsys@color@cmyk@stroke{0}{0.87}{0.68}{0.32}\pgfsys@color@cmyk@fill{0}{0.87}{0.68}{0.32}c_{1}=7(e-1)\sqrt{2}}$ concludes the proof. ∎
@@
On event $\mathcal{E}_{1}$, for all intervals $[s_{1},s_{2}]$, all depths $d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t}$, and all $B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$,
$\left|\sum_{t=s_{1}}^{s_{2}}\hat{\delta}_{t}(B^{\prime},B)-\sum_{t=s_{1}}^{s_{2}}\delta_{t}(B^{\prime},B)\right|\leq\underbrace{c_{1}\log(T)\sqrt{(s_{2}-s_{1})2^{d}\vee 4^{d}}}_{\text{Concentration}}+\underbrace{\frac{4(s_{2}-s_{1})}{2^{d}}}_{\text{Bias}}.$
Proof of Corollary 2.
This follows immediately by combining the concentration event (Proposition 2) with the bias control result from Proposition 3. ∎
@@ D.3 Properties of the eviction scheme of MDBE
On event $\mathcal{E}_{1}$, if a bin $B\in\mathcal{T}_{d}$ is evicted at some round $t\geq\tau_{l,m}$, then for all arm $x\in B$, there exists an interval $[s_{1},s_{2}]$ with $s_{1}\geq\tau_{l,m}$ and $s_{2}=t$ such that $x$ has incurred significant regret on this interval.
Proof of Proposition 4.
Assume $B\in\mathcal{T}_{d}$ has been evicted at a round $t$. By the design of MDBE, there exists an interval $[s_{1},s_{2}]$ with $s_{1}\geq\tau_{l,m}$ and $s_{2}=t$, such that
$\exists d^{\prime}\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t},\,\exists B_{\mathrm{parent}}\in\mathcal{B}_{[s_{1},s_{2}]}(d^{\prime})\text{ s.t. }B\subseteq B_{\mathrm{parent}}\text{ and }B_{\mathrm{parent}}\text{ is evicted at round }t\,,$
@@
Conditionally on concentration event $\mathcal{E}_{1}$, we have by Corollary 2, on interval $[s_{1},s_{2}]$,
@@
and therefore by eviction criteria ($\star$3) stated above
@@
for $c_{0}$ satisfying $c_{0}\geq 3+c_{1}$. Since $\sum_{t=s_{1}}^{s_{2}}\delta_{t}(B^{\prime},B_{\mathrm{parent}})\leq\sum_{t=s_{1}}^{s_{2}}\delta_{t}(B_{\mathrm{parent}})$, this implies that $B_{\mathrm{parent}}$ incurs significant regret on $[s_{1},s_{2}]$ (Definition 4). Since $x\in B_{\mathrm{parent}}$, it means that $x$ also incurred significant regret (Definition 1) by Proposition 1. ∎
@@
On event $\mathcal{E}_{1}$, $\forall\,[s_{1},s_{2}]\,,\forall d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t}\,,\,\forall B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$, if both $B$ and $B^{\prime}$ are not evicted during this interval, then
@@
where $c_{0}$ and $c_{1}$ are the positive numerical constants defined respectively in (italic-$\star$3) and Corollary 2.
Proof of Proposition 5.
If $B$ and $B^{\prime}$ are not evicted on $[s_{1},s_{2}]$, then for all $B^{\prime\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$, the eviction condition does not hold, i.e.
@@
By Corollary 2, the true cumulative gap satisfies
@@ D.4 Relating episode and significant shifts
We show that, with high probability, a new episode begins only if a significant shift occurs within it. This ensures a correspondence between episodes $[t_{l},t_{l+1}[$ and significant phases $[\tau_{i},\tau_{i+1}[$. In particular, each significant phase overlaps with at most two episodes.
@@
On event $\mathcal{E}_{1}$, for each episode $[t_{l},t_{l+1}[$, there exists at least one significant shift $\tau_{i}\in[t_{l},t_{l+1}[$.
Proof of Proposition 6.
Let $[\tau_{l,m},\tau_{l,m+1}[$ be the block during which episode $[t_{l},t_{l+1}[$ terminates. By the design of MDBE, the episode ends whenever $\mathcal{B}_{\mathrm{MASTER}}(m)=\emptyset$ at the end of this block. Thus, by the eviction condition ($\star$3), we have:
[…]
@@
Since we are on event $\mathcal{E}_{1}$, by Corollary 2:
@@
where we used $B(x)$ to denote the bin containing arm $x$.
This implies that bin $B(x)$ incurs significant regret on interval $[s_{1},s_{2}]$ (by Definition 4). Then, by Proposition 1, the arm $x\in B$ also incurs significant regret on $[s_{1},s_{2}]$ (Definition 1).
@@ D.5 Regret decomposition and discretion within one block
For each significant phase $[\tau_{i},\tau_{i+1}[$, we define $x_{i}^{\sharp}$ as the last safe arm (Definition 1), with ties broken arbitrarily. For all $t\in[\tau_{i},\tau_{i+1}[$, we define $x_{t}^{\sharp}=x_{i}^{\sharp}$.
@@
$\mathbb{E}\left[R(\pi,T)\right]=\sum_{t=1}^{T}\sup_{x\in\mathcal{X}}\mu_{t}(x)-\mathbb{E}\left[\sum_{t=1}^{T}\mu_{t}(x_{t})\right]=\mathbb{E}\left[\sum_{t=1}^{T}\delta_{t}(x_{t})\right]\,,$
@@
The difficulty lies in upper bounding the second sum $\mathbb{E}\left[\sum_{t}\delta_{t}(x_{t}^{\sharp},x_{t})\right]$. W.l.o.g., there are $T$ total episodes, and by convention we set $t_{l}=T+1$ if only $l-1$ episodes occur by round $t$. Summing over episodes gives
@@
We further decompose each episode $[t_{l},t_{l+1}[$ into blocks. Let $M_{l}=\lceil\log_{8}(7(t_{l+1}-t_{l})+1)-1\rceil$ be the maximum number of blocks. For notational convenience, extend $\tau_{l,m}=t_{l+1}-1$ for $m>M_{l}$. Then,
@@
Discretization trick within each block. Within each block $[\tau_{l,m},\tau_{l,m+1}[$, we relate the regret $\delta_{t}(x_{t}^{\sharp},x_{t})$ to the regret of their corresponding parent bins at depth $m$, introducing a discretization bias of order $1/2^{m}$:
@@
The first term captures the cumulative discretization error over one block $[\tau_{l,m},\tau_{l,m+1}[$ at depth $m$, and by Assumption 1, it is upper bounded as
@@
For the second term, by design of our sampling scheme (Algorithm 2), if $B_{t,m}$ is active, then conditionally on $B_{t,m}=B$ we have $x_{t}\sim\mathcal{U}(B)$. Otherwise, $x_{t}$ is sampled uniformly from one of its active parent at higher depth (that has no active child), and conditionally on $B_{t,m}=B$ we also have $x_{t}\sim\mathcal{U}(B)$, and therefore in both cases, $\forall t,\,\mathbb{E}\left[\mu_{t}(x_{t})\,|\,B_{t,m}=B\right]=\mu_{t}(B)$. This yields
@@
$\mathbb{E}\left[R(\pi,T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+6\sum_{l=1}^{T}\mathbb{E}\left[(t_{l+1}-t_{l})^{2/3}\right]$
@@
Using the fact that event $\mathcal{E}_{1}$ holds with probability at least $1-1/T^{3}$ (Proposition 2), and the fact that rewards are bounded in $[0,1]$, we get
$\mathbb{E}\left[R(\pi,T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+6\sum_{l=1}^{T}\mathbb{E}\left[\mathds{1}\!\left\{\mathcal{E}_{1}\right\}(t_{l+1}-t_{l})^{2/3}\right]+6\frac{T^{2}}{T^{3}}$
@@
We denote by $B^{\rm{last}}_{l,m}\in\mathcal{T}_{m}$ the last bin (with ties broken arbitrarily) at depth $m$ within block $[\tau_{l,m},\tau_{l,m+1}[$ that belongs to $\mathcal{B}_{\mathrm{MASTER}}(m)$.
@@
  • Replays within block $[\tau_{l,m},\tau_{l,m+1}[$ are scheduled at $\tau_{l,m}$, independently of observations prior to that round. Thus, for $t\in[\tau_{l,m},\tau_{l,m+1}[$ and $d<m$, $R_{t,d}$ is $\mathcal{F}_{\tau_{l,m}}$-measurable.

### v2

@@ Appendix D Proof of the upper bound of MDBE (Theorem˜2)
$\texttt{Replay}(s,d)$ denotes a replay at depth $d$ starting at round $s$.
@@ D.1 Bias of the mean estimates
$\forall[s_{1},s_{2}],\,\forall d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t},\,\forall B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]},\qquad\left|\sum_{t=s_{1}}^{s_{2}}\mathbb{E}_{t-1}\left[\hat{\delta}_{t}(B^{\prime},B)\right]-\sum_{t=s_{1}}^{s_{2}}\delta_{t}(B^{\prime},B)\right|\leq\frac{4(s_{2}-s_{1})}{2^{d}}\,.$
Proof of Proposition˜3.
@@
From Assumption˜1, for all bin $B\in\mathcal{T}_{d}$ and any $x\in B$, we have:
@@
Summing this inequality over the interval $s_{1},s_{2}]$ establishes the second claim. ∎
@@ D.2 Concentration of the mean estimates
Our sampling scheme (Algorithm˜2) is carefully constructed to ensure that the estimated gap between any two active bins at any active depth concentrates around its conditional expectation, especially by controlling its cumulative variance. Our result relies on this following inequality.
Lemma 1 (Beygelzimer et al. [2011]).
@@
Let $\mathcal{E}_{1}$ be the following event: for all intervals $s_{1},s_{2}]$, depths $d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t}$, and bins $B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$, we have
@@
$\geq\frac{1}{|\mathcal{B}_{t}(d^{\prime})\cap\mathrm{Children}(B_{d},d^{\prime})|}\frac{1}{2^{d}}$
$\geq\frac{2^{d}}{2^{d^{\prime}}}\frac{1}{2^{d}}$
$=\frac{1}{2^{d^{\prime}}}\,.$
Bound on each $M_{t}$. Recall the definition of $\hat{\delta}_{t}(B^{\prime},B)$,
@@
$\mathbb{E}_{t-1}\left[\hat{\delta}_{t}^{2}(B^{\prime},B)\right]\leq\frac{2}{(1/2^{d})^{2}}\cdot\frac{1}{2^{d}}=2^{d+1},\,$
where we use the fact that the probability that $x_{t}$ belong to $B_{d}$ or $B_{d}^{\prime}$ is upper bounded by $1/2^{d}$.
Final concentration bound. Applying Lemma˜1 over any interval $s_{1},s_{2}]$ for a fixed depth $d$ and for any pair of bins $B,B^{\prime}\in\mathcal{T}_{d}$, we get that with probability at least $1-\delta$,
@@
Choosing $\delta=1/T^{7}$ and $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}c_{1}=7(e-1)\sqrt{2}$} concludes the proof. ∎
@@
On event $\mathcal{E}_{1}$, for all intervals $s_{1},s_{2}]$, all depths $d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t}$, and all $B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$,
$\left|\sum_{t=s_{1}}^{s_{2}}\hat{\delta}_{t}(B^{\prime},B)-\sum_{t=s_{1}}^{s_{2}}\delta_{t}(B^{\prime},B)\right|\leq\underbrace{c_{1}\log(T)\sqrt{(s_{2}-s_{1})2^{d}\vee 4^{d}}}_{\text{Concentration}}+\underbrace{4(s_{2}-s_{1})/2^{d}}_{\text{Bias}}.$
Proof of Corollary˜2.
This follows immediately by combining the concentration event (Proposition˜2) with the bias control result from Proposition˜3. ∎
@@ D.3 Properties of the eviction scheme of MDBE
On event $\mathcal{E}_{1}$, if a bin $B\in\mathcal{T}_{d}$ is evicted at some round $t\geq\tau_{l,m}$, then for all arm $x\in B$, there exists an interval $s_{1},s_{2}]$ with $s_{1}\geq\tau_{l,m}$ and $s_{2}=t$ such that $x$ has incurred significant regret on this interval.
Proof of Proposition˜4.
Assume $B\in\mathcal{T}_{d}$ has been evicted at a round $t$. By design of MDBE, there exists an interval $s_{1},s_{2}]$ with $s_{1}\geq\tau_{l,m}$ and $s_{2}=t$, such that
[…]
@@
On the concentration event $\mathcal{E}_{1}$, we have by Corollary˜2, on interval $s_{1},s_{2}]$,
@@
and therefore by eviction criteria ($\star$3) we have
@@
for $c_{0}$ satisfying $c_{0}\geq 3+c_{1}$. Since $\sum_{t=s_{1}}^{s_{2}}\delta_{t}(B^{\prime},B_{\mathrm{parent}})\leq\sum_{t=s_{1}}^{s_{2}}\delta_{t}(B_{\mathrm{parent}})$, this implies that $B_{\mathrm{parent}}$ incurs significant regret on $s_{1},s_{2}]$ (Definition˜4). Since $x\in B_{\mathrm{parent}}$, it also implies that $x$ incurs significant regret (Definition˜1) by Proposition˜1. ∎
@@
On concentration event $\mathcal{E}_{1}$, $\forall\,[s_{1},s_{2}]\,,\forall d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t}\,,\,\forall B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$, if both $B$ and $B^{\prime}$ are not evicted during this interval, then
@@
where $c_{0}$ and $c_{1}$ are the positive numerical constants defined respectively in ($\star$3) and Corollary˜2.
Proof of Proposition˜5.
If $B$ and $B^{\prime}$ are not evicted on $s_{1},s_{2}]$, then for all $B^{\prime\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$, the eviction condition does not hold, i.e. we have
@@
By Corollary˜2, the true cumulative gap satisfies
@@ D.4 Relating episode and significant shifts
We show that, with high probability, a new episode begins only if a significant shift occurs within it. This ensures a correspondence between episodes $t_{l},t_{l+1}[$ and significant phases $\tau_{i},\tau_{i+1}[$. In particular, each significant phase overlaps with at most two episodes.
@@
On concentration event $\mathcal{E}_{1}$, for each episode $t_{l},t_{l+1}[$, there exists at least one significant shift $\tau_{i}\in[t_{l},t_{l+1}[$.
Proof of Proposition˜6.
Let $\tau_{l,m},\tau_{l,m+1}[$ be the block during which episode $t_{l},t_{l+1}[$ terminates. By the design of MDBE, the episode ends whenever $\mathcal{B}_{\mathrm{MASTER}}(m)=\emptyset$ at the end of this block. Thus, by the eviction condition ($\star$3), we have:
$\forall x\in\mathcal{X},\;\exists[s_{1},s_{2}]\subseteq[\tau_{l,m},\tau_{l,m+1}[,\;\exists d\in\bigcap_{s=s_{1}}^{s_{2}}\mathcal{D}_{s},\;\exists B\in\mathcal{B}_{[s_{1},s_{2}]}(d)\text{ with }x\in B,\;\exists B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)\;\text{s.t.}$
@@
Since we are on event $\mathcal{E}_{1}$, by Corollary˜2,
@@
where we used $B_{d}(x)$ to denote the bin at depth $d$ containing arm $x$.
This implies that bin $B_{d}(x)$ incurs significant regret on interval $s_{1},s_{2}]$ (by Definition˜4). Then, by Proposition˜1, the arm $x\in B$ also incurs significant regret on $s_{1},s_{2}]$ (Definition˜1).
@@ D.5 Regret decomposition and discretion within one block
For each significant phase $\tau_{i},\tau_{i+1}[$, we define $x_{i}^{\sharp}$ as the last safe arm (Definition˜1), with ties broken arbitrarily. For all $t\in[\tau_{i},\tau_{i+1}[$, we define $x_{t}^{\sharp}=x_{i}^{\sharp}$.
@@
$\mathbb{E}\left[R(\pi_{\texttt{MDBE}},T)\right]=\sum_{t=1}^{T}\sup_{x\in\mathcal{X}}\mu_{t}(x)-\mathbb{E}\left[\sum_{t=1}^{T}\mu_{t}(x_{t})\right]=\mathbb{E}\left[\sum_{t=1}^{T}\delta_{t}(x_{t})\right]\,$,
@@
The difficulty lies in upper bounding the second sum $\mathbb{E}\left[\sum_{t}\delta_{t}(x_{t}^{\sharp},x_{t})\right]$. Without loss of generality, there are $T$ total episodes, and by convention we set $t_{l}=T+1$ if only l1l-1 episodes occur by round $t$. Summing over episodes gives
@@
We further decompose each episode $t_{l},t_{l+1}[$ into blocks. Let $M_{l}=\lceil\log_{8}(7(t_{l+1}-t_{l})+1)-1\rceil$ be the maximum number of block within this episode. For notational convenience, we can extend $\tau_{l,m}=t_{l+1}-1$ for $m>M_{l}$. Then,
@@
Discretization trick within each block. Within each block $\tau_{l,m},\tau_{l,m+1}[$, we relate the regret $\delta_{t}(x_{t}^{\sharp},x_{t})$ to the regret of their corresponding parent bins at depth $m$, introducing a discretization bias of order $1/2^{m}$:
@@
The first term captures the cumulative discretization error over one block $\tau_{l,m},\tau_{l,m+1}[$ at depth $m$, and by Assumption˜1, it is upper bounded as
@@
For the second term, by design of our sampling scheme (Algorithm˜2), if $B_{t,m}$ is active, then conditionally on $B_{t,m}=B$ we have $x_{t}\sim\mathcal{U}(B)$. Otherwise, $x_{t}$ is sampled uniformly from one of its active parent at higher depth (that has no active child), and conditionally on $B_{t,m}=B$ we also have $x_{t}\sim\mathcal{U}(B)$, and therefore in both cases, $\forall t,\,\mathbb{E}\left[\mu_{t}(x_{t})\,|\,B_{t,m}=B\right]=\mu_{t}(B)$. This yields
@@
$\mathbb{E}\left[R(\pi_{\texttt{MDBE}},T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+6\sum_{l=1}^{T}\mathbb{E}\left[(t_{l+1}-t_{l})^{2/3}\right]$
@@
Using the fact that event $\mathcal{E}_{1}$ holds with probability at least $1-1/T^{3}$ (Proposition˜2), and the fact that rewards are bounded in $[0,1]$, we get
$\mathbb{E}\left[R(\pi_{\texttt{MDBE}},T)\right]$ $\leq\log(T)\sum_{i=0}^{\tilde{L}_{T}}(\tau_{i+1}-\tau_{i})^{2/3}+6\sum_{l=1}^{T}\mathbb{E}\left[\mathds{1}\!\left\{\mathcal{E}_{1}\right\}(t_{l+1}-t_{l})^{2/3}\right]+6\frac{T^{2}}{T^{3}}$
@@
We denote by $B^{\rm{last}}_{l,m}\in\mathcal{T}_{m}$ the last bin (with ties broken arbitrarily) at depth $m$ within block $\tau_{l,m},\tau_{l,m+1}[$ that belongs to $\mathcal{B}_{\mathrm{MASTER}}(m)$.
@@
  • Replays within block $\tau_{l,m},\tau_{l,m+1}[$ are scheduled at $\tau_{l,m}$, independently of observations prior to that round. Thus, for $t\in[\tau_{l,m},\tau_{l,m+1}[$ and $d<m$, $R_{t,d}$ is $\mathcal{F}_{\tau_{l,m}}$-measurable.